# Wanda++: Pruning Large Language Models via Regional Gradients

### v2

@@ Abstract
Large Language Models (LLMs) pruning seeks to remove unimportant weights for inference speedup with minimal performance impact. However, existing methods often suffer from performance loss without full-model sparsity-aware fine-tuning. This paper presents Wanda++, a novel pruning framework that outperforms the state-of-the-art methods by utilizing decoder-block-level regional gradients. Specifically, Wanda++ improves the pruning score with regional gradients for the first time and proposes an efficient regional optimization method to minimize pruning-induced output discrepancies between the dense and sparse decoder output. Notably, Wanda++ improves perplexity by up to 32% over Wanda in the language modeling task and generalizes effectively to downstream tasks. Further experiments indicate our proposed method is orthogonal to sparsity-aware fine-tuning, where Wanda++ can be combined with LoRA fine-tuning to achieve a similar perplexity improvement as the Wanda method. The proposed method is lightweight, pruning a 7B LLaMA model in under 10 minutes on a single NVIDIA H100 GPU. [^1] The unofficial implementation of the Wanda++ method, contributed by the community, is available at https://github.com/TTTTTTris/wandaplus.
@@ 1 Introduction
The growing size of Large Language Models (LLMs) improves performance (Devlin et al., 2018; Touvron et al., 2023) at the cost of memory consumption and inference latency. For example, hosting an LLaMA-2-70B model needs at least four A100-40GB GPUs with the time to first token (TTFT) exceeding 100 milliseconds (Agarwal, 2023). To address these challenges, various model compression approaches, including weight decomposition (Hsu et al., 2022; Yang et al., 2024), quantization (Lin et al., 2024; Tian et al., 2023), and pruning (Sun et al., 2023; Xu et al., 2024), have been explored. Among pruning methods, post-training LLM pruning approaches, such as SparseGPT (Sun et al., 2023) and Wanda (Sun et al., 2023), have gained attention as they circumvent the prohibitive memory overhead associated with traditional in-training pruning techniques (Han et al., 2015; Frankle & Carbin, 2018). However, post-training pruning often leads to substantial performance degradation, limiting its effectiveness for efficient LLM deployment.
To mitigate the performance degradation, GBLM and Pruner-Zero (Das et al., 2023; Dong et al., 2024) propose improved pruning criteria that enhance the layer-wise Wanda score by incorporating gradient information obtained through full-model backpropagation. Meanwhile, other approaches focus on recovering model performance through sparsity-aware tuning (Sun et al., 2023) or distillation (Liang et al., 2023). Although these methods effectively reduce pruning-induced degradation, they suffer from impractical memory requirements and excessive pruning time due to the high computational cost of full-model backpropagation. This raises an important question:
@@
In this paper, we propose Wanda++ pruning framework to leverage gradients at the “decoder-block" level, termed regional gradients, which shows significant improvement compared to Wanda Sun et al. (2023) without largely sacrificing the pruning efficiency. Compared to the Wanda method, the proposed approach efficiently incorporates crucial gradient information during pruning, which has been shown to provide non-trivial insights in prior works such as Optimal Brain Surgeon (Hassibi et al., 1993) and GBLM (Das et al., 2023). Notably, the regional gradient can be computed by loading and processing backpropagation for only a single decoder block at a time, making gradient calculation feasible regardless of the LLM’s size.
Our proposed pruning framework consists of two components that leverage regional gradients: Regional Gradient Score (RGS) and Regional Optimizer (RO). The RGS demonstrates the effectiveness of replacing the full-model gradient used in the GBLM method with a regional gradient, computed via backpropagation on a loss function defined as the $\ell_{2}$ norm of the decoder output. Meanwhile, the RO method performs lightweight, decoder-block-level weight updates after each pruning step, mitigating pruning-induced loss by adjusting weights based on a small calibration dataset. This is achieved by considering the difference between the outputs of dense and pruned decoder blocks, ensuring efficient local recovery. Our contributions can be summarized as follows:
@@
  • We propose Wanda++, a lightweight yet effective framework that prunes LLMs using regional gradients, achieving significant performance improvements without requiring full-model backward.
@@
  • As shown in Fig. 1, Wanda++ effectively mitigates pruning-induced degradation in a non-incremental manner and generalizes well to zero-shot downstream tasks.
@@
  • The proposed method is orthogonal to previous sparsity-aware fine-tuning approaches and achieves a similar perplexity improvement as Wanda when combined with LoRA fine-tuning.
@@ 2 The Wanda++ Framework
In this section, we introduce Wanda++, a framework designed to efficiently reduce pruning-induced degradation through a two-stage process. The algorithm flow is summarized in Algorithm 1, which iteratively performs pruning based on the regional pruning score followed by rapid weight updates using the RO process. Finally, a last pruning step is performed to obtain the pruned weight for each decoder blocks.
@@ 2.1 Regional Gradient Score
As the first stage of Wanda++, we obtain the Regional Gradient Score (RGS) for in-block pruning. We start with constructing an RGS loss function for obtaining the gradient of each weight matrix. Given a model with $L$ decoder blocks, we represent the set of input hidden states for the $l$-th decoder block specifically as $\mathcal{X}^{l}=\{\bm{X}_{1}^{l},\cdots,\bm{X}_{N}^{l}\}$ and define the decoder block function as $f^{l}(\bm{X}^{l}_{n})$ with input $\bm{X}^{l}_{n}\in\mathcal{X}^{l}$. The RGS loss for $l$-th block is defined as $\mathcal{L}_{RGS}^{l}(\bm{X}^{l}_{n})=\|f^{l}(\bm{X}^{l}_{n})\|_{2}$. By performing a single backpropagation through a certain decoder block regarding the $\mathcal{L}^{l}$, we can efficiently obtain the regional gradient as $\nabla_{\bm{W}_{ij}}\mathcal{L}_{RGS}^{l}(\bm{X}^{l}_{n})$.
We compute the regional gradient $\nabla\mathcal{L}_{RGS}^{l}(\bm{X}^{l}_{n})$ only once during the iteratively pruning and weights update process. Here, we replace the full-model gradient utilized in GBLM score with our regional gradient to obtained the RGS score. Note that both RGS and GBLM score only compute the gradient once. Thus, to further capture the interdependence of layers within the same block, the Wanda score is integrated, which tracks how pruning a single linear layer affects other layers in the block. In conclusion, ourRGS criterion is summarized as follows:
$\bm{S}_{ij}=\left(\frac{\alpha}{N}\textstyle{\sqrt{\sum_{n=1}^{N}(\nabla\mathcal{L}_{RGS}^{l}(\mathbf{X}_{n}^{l})^{2})_{ij}}}+\|\mathbf{X}_{j}\|_{2}\right)\cdot|\bm{W}_{ij}|$ (1)
where $N$ represents the total number of input samples used for pruning and the scaling factor $\alpha$ is a constant to balance the magnitude of the gradient and input activation terms. We choose the value of $\alpha$ to be 100 based on our ablation study in Appendix D.
@@ 3 Experiment
We closely follow the experimental setup of previous work like Wanda Sun et al. (2023) on unstructured sparsity, 2:4 sparsity, and 4:8 sparsity. Regarding the model, we consider OpenLLaMA (3B/7B/70B), LLaMA-1 (7B/13B/30B/65B) and LLaMA-3.1 (8B). By default, we randomly select 128 samples from the C4 training data for regional optimization and evaluate perplexity on both the C4/Wikitext test datasets. For zero-shot evaluation, we use the Harness evaluation toolkit (Gao et al., 2024). We test Wanda++ RGS for the performance of RGS criteria itself. Further experiments on the sensitivity analysis regarding sample sizes , latency reduction, and the ablation studies of the RO component can be found in Appendix A, B, E. All experiments are conducted on the H100 GPU.
Perplexity: We report our results in Table 1. We consider SparseGPT, Wanda and GBLM as baselines to compare with our proposed method. In all experiments, our method consistently shows superior perplexity compared to the baseline methods. For the OpenLLaMA-3B, the relative perplexity reductions are 32.1% and 25.5% for 2:4 and 4:8 sparsity, respectively. On LLaMA-1-7B model with 2:4 sparsity, the relative reductions are 19%, 20%, 7%, and 11% for model sizes 7B, 13B, 30B, and 65B, respectively. Compared to the GBLM score, the Wanda++ RGS achieves similar performance, demonstrating that regional gradients serve as a strong replacement for full-model gradients, especially for the 30B model. We were unable to obtain results for the 65B/70B models using the GBLM and SparseGPT methods due to computational resource limitations.
@@
Table 1: WikiText perplexity comparison across baselines. Bold highlights relative perplexity improvements over Wanda of 5% or more.
Zero-Shot Accuracy: We compare the downstream performance with Wanda. Although no consistent pattern shows, Wanda++ in general yields the best performance among the three pruning methods including Wanda which leads in BoolQ task. Compared to the margin from perplexity evaluation, the improvement from Wanda++ against Wanda++ RGS is less salient. This is reasonable as RO is conducted on C4 dataset without optimizing any downstream tasks. Note that for Mrpc and RTE tasks, Wanda++ outperforms Wanda by 46% and 24%, close to the accuracy of the dense baseline.
Method Wic Mrpc Hellaswag Arc_easy Arc_challenge Winogrande BoolQ RTE OBQA Mean
Baseline 49.84 69.12 56.96 75.29 41.80 70.00 75.02 66.43 34.40 59.87
Wanda 48.75 46.81 41.66 59.34 27.47 61.96 69.60 49.82 23.80 47.69
Wanda++ RGS 49.37 (1%) 64.46 (38%) 41.43 (-1%) 62.42 (5%) 31.06 (13%) 62.83 (1%) 67.95 (-2%) 58.48 (17%) 23.40 (-2%) 51.27
Wanda++ 50.00 (2%) 68.38 (46%) 45.31 (8%) 63.72 (7%) 29.27 (6%) 65.04 (4%) 67.80 (-2%) 62.09 (24%) 24.80 (4%) 52.93

### v3

@@ Abstract
Large Language Models (LLMs) pruning seeks to remove unimportant weights for inference speedup with minimal accuracy impact. However, existing methods often suffer from accuracy degradation without full-model sparsity-aware fine-tuning. This paper presents Wanda++, a novel pruning framework that outperforms the state-of-the-art methods by utilizing decoder-block-level regional gradients. Specifically, Wanda++ improves the pruning score with regional gradients for the first time and proposes an efficient regional optimization method to minimize pruning-induced output discrepancies between the dense and sparse decoder output. Notably, Wanda++ improves perplexity by up to 32% over Wanda in the language modeling task and generalizes effectively to downstream tasks. Moreover, despite updating weights with regional optimization, Wanda++ remains orthogonal to sparsity-aware fine-tuning, further reducing perplexity with LoRA in great extend. Our approach is lightweight, pruning a 7B LLaMA model in under 10 minutes on a single H100 GPU. [^1] The official code is unavailable due to legal constraints. However, an unofficial reproduction of the Wanda++ method, contributed by the community, is available at https://github.com/TTTTTTris/wandaplus.
Wanda++: Pruning Large Language Models via Regional Gradients
[…]
@@ 1 Introduction
The growing size of Large Language Models (LLMs) improves performance (Touvron et al., 2023; Intelligence, 2024a, b) at the cost of memory consumption and inference latency. For example, loading the weights of LLaMA-2 70B requires 140 GB of GPU memory in FP16 format. Hosting a such model even with a batch size of 1 and 512 input tokens needs at least four A100-40GB GPUs, with the time to first token (TTFT) exceeding 100 milliseconds (Agarwal, 2023). To address these challenges, various model compression approaches, including weight decomposition (Hsu et al., 2022; Yang et al., 2024; Ghiasvand et al., 2024), quantization (Lin et al., 2024; Zhou et al., 2025), and pruning (Sun et al., 2023; Frantar and Alistarh, 2023; Zhang et al., 2024), have been explored. Similar to LLM quantization, many recent LLM pruning methods have shifted from in-training approaches (Han et al., 2015b; Frankle and Carbin, 2018) to post-training methods, as seen in SparseGPT (Frantar and Alistarh, 2023) and Wanda (Sun et al., 2023). However, unlike post-training quantization like AWQ (Lin et al., 2024), which compresses model weights almost losslessly by 4$\times$ (from 16-bit to 4-bit), these pruning methods have yet to achieve comparable levels of performance.
@@
To mitigate the performance degradation, GBLM and Pruner-Zero (Das et al., 2023; Dong et al., 2024) propose improved pruning criteria that enhance the layer-wise Wanda score by incorporating gradient information obtained through full-model backpropagation (BP). These studies highlight the importance of gradient information, demonstrating that it provides valuable insights even though Wanda assumes that the gradients of a fully trained network are small and contribute little when higher-order terms are considered. Meanwhile, other approaches focus on recovering model performance through full-model sparsity-aware tuning Sun et al. (2023) or distillation (Liang et al., 2023). However, all these methods, which heavily depend on full-model loss, suffer from impractical memory requirements and excessive pruning time due to the high computational cost of full-model backpropagation. This raises the question:
@@
In this paper, we propose Wanda++ pruning framework to leverage gradients at the “decoder-block" level, termed regional gradients, which shows significant improvement compared to Wanda Sun et al. (2023) without sacrificing the pruning efficiency. Compared with traditional full model BP, the regional gradient can be obtained by only loading and computing gradients per single decoder block, which largely reduce the GPU hours required. Compare with previous gradient-free method like Wanda, the regional gradient can improve the pruning performance by accessing the gradient information. Wanda++ prune each decoder block by iteratively prune the model based on a Regional Gradient Score (RGS) and slightly update the weights with our proposed Regional Optimizer (RO).
To effectively obtain the regional gradient in the RGS score, we compute the regional gradient through a backward process on a regional loss, which is defined as the $\ell_{2}$ norm of each decoder’s output hidden states. To involve the regional gradient into the pruning score, we follow the design of GBLM score Das et al. (2023). Note that the regional gradient is only computed once during the iterative pruning process of each decoder block and reused for all RO iterations to further reduce the computation. As a result, the regional gradient may become less accurate after weights being updated during the RO process. The design of GBLM score provide an effective way to reflect changes by fetch the input per in-block layer and blend it into the RGS along with regional gradients.
For the RO in Wanda++, we construct a simple loss function between the outputs of the dense and pruned decoder blocks to update the model weights within each block. The concept of regional optimization traces back to previous local optimization approaches in convolutional neural networks (CNNs), which aimed to optimize models by focusing on individual convolutional layers (Wang et al., 2021). Differently, RO in Wanda++ extends this idea for mitigating pruning-induced loss on a small portion of calibration data at each LLM decoder-block, instead of optimizing the cross-entropy loss for classification tasks.
Our proposed framework achieves up to a 32% reduction in WikiText perplexity at 2:4 sparsity, compared to the Wanda method under the same experimental setup, as shown in Figure. 1. Our contributions can be summarized as follows:
@@
  • We propose Wanda++, a lightweight yet effective framework that prunes LLMs using regional gradients, achieving performance improvements without requiring access to full model gradients.
@@
  • Wanda++ effectively lowers the pruning-induced degradation in a non-incremental way and generalize well in zero-shot downstream tasks.
@@
  • The RO method in Wanda++ is orthogonal to previous full-model sparsity-aware fine-tuning methods and has been shown to achieve a similar perplexity improvement as Wanda after applying LoRA fine-tuning.
2 Related Work
Network Pruning: The concept of pruning neural networks has been explored for decades, beginning with foundational works such as (LeCun et al., 1989; Hassibi et al., 1993). In addition to widely studied unstructured pruning approaches, structured pruning methods, which remove entire subnets of a network or rows/columns within weight matrices, are more easily supported on hardware for inference speedup (Liu et al., 2017; Shen et al., 2022). Research in this area has focused on analyzing input/output activation statistics to identify the most suitable neurons for pruning (Bai et al., 2021; Molchanov et al., 2022). However, even at 50% sparsity, structured pruning often results in non-trivial performance degradation (Ashkboos et al., 2024). Beyond these methods, semi-structured pruning (Pool et al., 2021; Fang et al., 2022, 2023), such as 2:4 sparsity, shows greater resilience at 50% sparsity and can effectively reduce runtime latency with throughput improvement on NVIDIA’s recent hardware. Therefore, this work primarily focuses on unstructured and semi-structured patterns, although the proposed Wanda++, RO in particular, is sparsity pattern agnostic.
LLM Pruning: As LLMs continue to grow in size, scaling traditional pruning methods to accommodate them presents significant challenges. Traditional pruning methods, which typically require full model retraining, demand substantial computational resources, making them impractical in the era of LLMs. A notable trend in LLM pruning is the adoption of post-training pruning methods (Frantar and Alistarh, 2023; Sun et al., 2023; Das et al., 2023; Zhang et al., 2025), which develop specific pruning scores to determine the importance of different weight elements. Additionally, SliceGPT (Ashkboos et al., 2024) focuses on structured pruning by slicing rows or columns of weight matrices based on input eigenvectors and eigenvalues.
Pruning with Gradient Information: The use of gradient information during the training process has been extensively studied in two ways. The first set of methods directly incorporates gradient information to refine their pruning scores at a finer granularity. Recent developments, such as GBLM and Pruner-Zero (Das et al., 2023; Dong et al., 2024) use gradients obtained through the full model backward with respect to cross entropy loss to achieve better performance, though they are time and memory infeasible for large-scale model.
The second set of methods focuses more on incorporating gradient information during sparsity-aware distillation or fine-tuning. For example, Liang et al. (2023) explores a task-agnostic distillation approach combined with iterative pruning to achieve strong performance on natural language understanding tasks. Additionally, full-model or LoRA fine-tuning has been considered as an auxiliary step in various prior pruning works that emphasize single-shot performance to further reduce the pruning degradation, such as Wanda Sun et al. (2023) and SliceGPT Ashkboos et al. (2024).
3 Prior Solutions
Existing solvers for post-training LLM pruning often start with magnitude pruning (Han et al., 2015b), which removes individual weights based on their magnitudes. This approach has been widely adopted as a baseline method in both vision (Han et al., 2015a) and language processing (Gale et al., 2019). Magnitude pruning is built on the assumption that neurons with larger weight elements contribute higher gradients when input features have similar magnitudes. However, this assumption does not always hold true for LLMs, where input features can differ significantly in scale, as observed in (Dettmers et al., 2022).
To address this challenge, Sun et al. propose the Wanda method (Sun et al., 2023), which introduces a pruning criterion that multiplies the weight magnitude by the input activation on an element-wise basis. Specifically, given a linear layer with a weight matrix $\bm{W}\in\mathbb{R}^{d_{out}\times d_{in}}$ and input $\bm{X}\in\mathbb{R}^{(B\times L)\times d_{in}}$ assuming the batch size is $B$, sequence length is $L$, the Wanda pruning score for the weight element connect input $j$ and output $i$ is defined as:
$\bm{S}_{ij}=|\bm{W}_{ij}|\cdot\|\bm{X}_{j}\|_{2},$ (1)
where $\|\bm{X}_{j}\|_{2}$ represent the L2 norm of the $j$-th input channel $\bm{X}_{j}$. The Wanda score is further improved by Das et al. (2023), which shows the first order gradient is still crucial even though Wanda score consider the higher-order terms and gives the GBLM score:
$\bm{S}_{ij}=(\alpha\bm{G}_{ij}+\|\bm{X}_{j}\|_{2})\cdot|\bm{W}_{ij}|,$ (2)
where $\bm{G}$ is full model gradient with respect to the cross-entropy loss between the prediction and labels. Next, we introduce how to involve the regional gradient into the pruning score and perform a effective and lightweight regional optimization.
Figure 2: Illustration of Wanda++, which leverages regional gradients in two ways: first, we estimate the block-level gradient $\mathbf{G}$, used in Regional-Gradient Scoring (RGS) for layer-wise pruning in each block. Next, we calculate the pruning-induced loss $\mathcal{L}{ro}$ for lightweight Regional Optimization (RO). RO can be iterated multiple times to infer better pruning masks for each decoder block.
@@ 4 The Wanda++ Framework
In this section, we use the Wanda method to highlight the drawbacks of the linearity assumption in previous layer-wise post-training pruning approaches. We then introduce our Wanda++ framework, which efficiently reduces pruning degradation with a two-stage process for each decoder. This involves layer-wise pruning within blocks based on a new pruning criterion called the Regional Gradient Score (RGS) and Regional Optimization (RO), as shown in Figure 2. Wanda++ is efficient as it operates at each decoder block where its RO requires only 5 iterations.
@@ 4.1 Regional Gradient Score
As the first stage of Wanda++, we obtain the Regional Gradient Score (RGS) for in-block layer-wise pruning. We start with constructing an RGS loss function for obtaining the gradient of each weight matrix. Given a model with $L$ decoder blocks, we represent the set of input hidden states for the $l$-th decoder block specifically as $\mathcal{X}^{l}=\{\bm{X}_{1}^{l},\cdots,\bm{X}_{N}^{l}\}$ and define the decoder block function as $f^{l}(\bm{X}^{l}_{n})$ with input $\bm{X}^{l}_{n}\in\mathcal{X}^{l}$. The RGS loss for $l$-th block is defined as $\mathcal{L}_{RGS}^{l}(\bm{X}^{l}_{n})=\|f^{l}(\bm{X}^{l}_{n})\|_{2}$. By performing a single BP through a certain decoder block regarding the regional loss $\mathcal{L}^{l}$, we can efficiently obtain the stochastic estimated gradient magnitude for each weight matrix by taking the absolute value for the estimated gradient $\nabla_{\bm{W}_{ij}}\mathcal{L}_{RGS}^{l}(\bm{X}^{l}_{n})$.
By performing the element-wise multiplication between the weight and estimated gradient magnitude, we have the initial regional gradient-based score for each weight element $W_{ij}$:
$\bm{S}_{ij}=(\sqrt{\frac{\sum^{N}_{n=1}\nabla\mathcal{L}_{RGS}^{l}(\bm{X}^{l}_{n})^{2}}{N}})_{ij}\cdot|\bm{W}_{ij}|,$ (3)
where $N$ represents the total number of input samples used for pruning. For simplicity, we record the score in Eq. (3) as $\bm{S}_{ij}=\bm{G}_{ij}\cdot|\bm{W}_{ij}|$.
To limit the overhead of computing the pruning criterion, we estimate the regional gradient (Eq. 4) only once per decoder block. This can be sub-optimal because pruning one layer in the $l$-th block affects the regional gradients of other layers. To address this, we blend in the layer-wise Wanda score, which tracks how pruning in one layer impacts the input of the next. Our RGS criterion is summarized as follows:
$\bm{S}_{ij}=(\alpha\bm{G}_{ij}+\|\bm{X}_{j}\|_{2})\cdot|\bm{W}_{ij}|,$ (4)
where the scaling factor $\alpha$ is a constant to balance the magnitude of the gradient and input activation terms. For simplicity, we adopt the same scaling value of 100 as in (Das et al., 2023) although it can be model specific (see the ablation study on $\alpha$ in Appendix B.2).
4.2 Regional Optimization
4.3 Algorithms
We summarize the Wanda++ algorithm flow in Alg. 1, as also shown in Figure 2. For each decoder block, we perform layer-wise RGS pruning (step 5) followed by a round of the RO process (steps 6-8) for $K$ iterations. An additional RGS pruning (step 11) is required to restore sparsity after RO. To detail how RGS is computed in steps 5 and 11, we provide PyTorch pseudo-code in Appendix A.
@@ 5 Experiment
We evaluate the proposed pruning method based on four criteria. The first is perplexity in next-token prediction, which is the main objective of LLM pre-training. A lower perplexity is often a strong indicator of superior performance in language understanding tasks. The second criterion is the performance on zero-shot/few-shot NLP tasks, such as text classification, question answering, and text generation. This is to ensure that the gradients involved in regional optimization do not lead to overfitting for downstream tasks. The third criterion is pruning time and memory consumption. Finally, we examine the actual latency reduction under both FP16 and FP8 quantization settings for various batch sizes and input/output lengths.
5.1 Experimental Setup
Regarding the model, we consider OpenLLaMA (3B/7B/70B), LLaMA-1 (7B/13B/30B/65B) and LLaMA-3.1 (8B). We follow the same settings as in (Sun et al., 2023) to examine pruning performance on unstructured sparsity, 2:4 sparsity, and 4:8 sparsity. By default, we randomly select 128 samples from the C4 training data for regional optimization and evaluate perplexity on both the C4 validation and Wikitext test datasets. For zero-shot evaluation, we use the Harness evaluation toolkit (Gao et al., 2024). All experiments are conducted on the NVIDIA H100 GPU, where a single GPU is sufficient for models with 13B or fewer parameters and 4 GPUs are used for 65B and 70B models. Latency and model size are measured via TensorRT-LLM. Regarding the proposed pruning methods, by default, Wanda++ enables both RGS and RO; Wanda++ RGS excludes RO, using only RGS for pruning, while Wanda++ RO applies the Wanda score (Sun et al., 2023) for pruning and weight updates within each decoder block.
@@ 5.1 Experimental Setup
Table 1: Wikitext perplexity comparison on LLaMA-1, OpenLLaMA, and LLaMA-3.1 model families. * indicates results from either the previous paper (Sun et al., 2023). - means results are not applicable from running their source code directly or out-of-memory issue. Bold highlights relative perplexity improvements over Wanda of 5% or more.
Figure 3: Perplexity on the C4 validation (left) and Wikitext test sets (right) as the sparsity ratio increases by gradually pruning two additional decoding blocks at a time, up to all 26 blocks. Wanda++ leverages regional gradients and significantly outperforms Wanda: our 2:4 results are comparable to Wanda’s 4:8 counterparts.
5.2 Perplexity
First, we consider OpenLLaMA-3B, a relatively small model, for a more contrastive perplexity comparison between Wanda and our proposed method. As shown in Figure 3, we begin by pruning the first 2 decoder blocks, then gradually apply N:M sparsity to two more decoder blocks until all 26 blocks are pruned. The LM head and embeddings are excluded from pruning. Perplexity results are reported on both the C4 validation dataset and the WikiText test dataset. When all decoder blocks are pruned with the N:M pattern, resulting in a 50% sparsity ratio, our method outperforms Wanda by a noticeable margin: on the C4 validation dataset, with 2:4 sparsity, perplexity is reduced from 36.5 to 23.9, a relative reduction of 34.4%; with 4:8 sparsity, perplexity is reduced from 23.1 to 16.4, a relative reduction of 29.0%. Similarly, on the WikiText test dataset, the relative perplexity reductions are 32.1% and 25.5% for 2:4 and 4:8 sparsity, respectively. The margin generally increases with the sparsity ratio, although in the lower sparsity ratio region, particularly on WikiText (right plot), the benefit is obscure. Note that, in higher sparsity ratios, our method can achieve comparable or superior perplexity results under 2:4 sparsity compared to Wanda for 4:8 sparsity, a much more relaxed sparsity setting.
A more comprehensive comparison of perplexity is summarized in Table 1, where LLaMA-1 models with four different sizes are included along with OpenLLaMA 7B and 70B. We consider SparseGPT and Wanda as two baseline pruning methods to compare with our method, Wanda++ w. RO. To understand the contribution of each enhancement in our method, we also report results from Wanda++ RO, where the Regional Optimizer (RO) is enabled for each decoder block, and Wanda++, where regional gradients are used along with the L2 norm of the input data for every layer in each decoder block. Regarding the sparsity patterns, we consider 50% unstructured pruning, 2:4, and 4:8 semi-structured pruning methods, although 2:4 sparsity is the most commonly supported in hardware for runtime acceleration.
Between Wanda++ RO and Wanda++ RGS, leveraging regional gradients within the decoder block-level optimization proves more effective in mitigating pruning-induced degradation. Furthermore, applying RO to a better pruning criterion, as shown in Wanda++ w. RO compared to Wanda++ RO, further improves performance. We compare our method (the strongest among those three in italics) to Wanda and report the relative perplexity reductions in Table 1. The most noticeable benefit is observed with 2:4 sparsity: on LLaMA-1, the relative reductions are 19%, 20%, 7%, and 11% for model sizes 7B, 13B, 30B, and 65B, respectively. Our method is most effective in reducing pruning degradation for smaller models, where higher pruning degradations usually occur, which is also the case for OpenLLaMA models. In all experiments, our method consistently shows superior perplexity compared to the baseline methods. However, for 4:8 sparsity and unstructured pruning, where perplexity values are generally lower than those with 2:4 sparsity, the benefits of our method become less salient. Nonetheless, the improvement remains more substantial compared to what Wanda achieves on top of SparseGPT.
Sparsity-aware model fine-tuning may further improve the performance. We further show our Wanda++ method is orthogonal to LoRA fine-tuning Hu et al. (2021) method in Section. 5.6. Also Another method distills smaller LLMs from a 15B pre-trained model but requires 16 DGX A100 nodes (Muralidharan et al., 2024) (128 80GB GPUs in total), whereas our method only needs one for the similar model size.
Method Wic Mrpc Hellaswag Arc_easy Arc_challenge Winogrande BoolQ RTE MMLU
Baseline 49.84 69.12 56.96 75.29 41.80 70.00 75.02 66.43 35.10
Wanda 48.75 46.81 41.66 59.34 27.47 61.96 69.60 49.82 25.85
GBLM 49.32 65.31 41.80 61.43 30.45 63.24 71.20 57.43 26.34
Wanda++ RGS 49.37 (1%) 64.46 (38%) 41.43 (-1%) 62.42 (5%) 31.06 (13%) 62.83 (1%) 67.95 (-2%) 58.48 (17%) 26.40 (-2%)
Wanda++ 50.00 (2%) 68.38 (46%) 45.31 (8%) 63.72 (7%) 29.27 (6%) 65.04 (4%) 67.80 (-2%) 62.09 (24%) 27.52 (6%)